Don runs a charity fruit sale, selling boxes of oranges for $11 and boxes of grapefruit for $10. If he sold a total of 762 boxes and is total income was $8125, how many boxes of oranges did he sell?
Answer by checkley71(8403)

10(762-x)+11X=8125 OR 7620-10x+11x=8125 or x=8125-7620 or x=505 boxes of oranges sold thus he also sold 762-505=257 boxes of grapefruit
505*11+257*10=8125 or 5555+2570=8125 or 8125=8125